# From Age Estimation to Age-Invariant Face Recognition: Generalized Age Feature Extraction Using Order-Enhanced Contrastive Learning

### v1

@@ Abstract
Generalized age feature extraction is crucial for age-related facial analysis tasks, such as age estimation and age-invariant face recognition (AIFR). Despite the recent successes of models in homogeneous-dataset experiments, their performance drops significantly in cross-dataset evaluations. Most of these models fail to extract generalized age features as they only attempt to map extracted features with training age labels directly without explicitly modeling the natural progression of aging. In this paper, we propose Order-Enhanced Contrastive Learning (OrdCon), which aims to extract generalized age features to minimize the domain gap across different datasets and scenarios. OrdCon aligns the direction vector of two features with either the natural aging direction or its reverse to effectively model the aging process. The method also leverages metric learning which is incorporated with a novel soft proxy matching loss to ensure that features are positioned around the center of each age cluster with minimum intra-class variance. We demonstrate that our proposed method achieves comparable results to state-of-the-art methods on various benchmark datasets in homogeneous-dataset evaluations for both age estimation and AIFR. In cross-dataset experiments, our method reduces the mean absolute error by about 1.38 in average for age estimation task and boosts the average accuracy for AIFR by 1.87%.
@@ I Introduction
Despite the importance of generalized age features, current age feature extraction methods mainly rely on conventional supervised learning, where models are designed to capture facial features associated with aging by directly mapping them to age labels [13]. A common approach involves using deep neural networks to automatically learn age features such as skin texture, wrinkles, and facial structure changes [14, 15]. While these methods have demonstrated success in controlled datasets and environments, they fall short of generalizing to diverse scenarios with unfamiliar variability in aging patterns across individuals [16].
To elaborate, one of the main challenges in age feature extraction is overfitting. Models trained with conventional supervised learning techniques optimize loss functions to closely match the training data, thus learning highly specific features tied to the age distributions and demographic characteristics in that data [17]. This hampers their ability to generalize to unseen data, particularly in diverse environments where aging processes vary significantly across individuals [18].
@@
Additionally, models trained with conventional supervised learning techniques often struggle to capture the inherent continuity and progression of age-related changes [19]. Most supervised approaches treat each sample independently, disregarding the sequential and gradual nature of the aging process [17, 14]. This inability to capture smooth transitions in facial features over time can result in inconsistent or unrealistic predictions that do not reflect the continuous nature of aging [20].
@@
Figure 1: Visual representation of the age features learned using the proposed method. Top: Feature space aligned with the direction of age progression. Bottom: Features from two individuals at different ages, following the learned age progression. Best view in color.
@@
To address the aforementioned limitations, we propose a novel method called Order-Enhanced Contrastive Learning (OrdCon). OrdCon leverages both order learning [21] and contrastive learning [22] to enhance the generalization of age feature extraction by explicitly modeling how features evolve with age progression. Unlike conventional contrastive learning that contrast image pairs [23], OrdCon contrasts the direction vector between features in the feature space. Specifically, it aligns the direction vectors from younger to older faces with the natural direction of age progression. This approach allows the model to capture continuous and gradual transitions in facial features, effectively addressing the lack of representation learning in conventional methods.
@@
Furthermore, the proposed method employs proxy-based metric learning [24] to reduce intra-class variance and create a structured feature space. In this approach, proxies are learnable entities that represent the centers of specific age clusters in the feature space. The model contrasts extracted features relative to these proxies, effectively minimizing intra-class variance by ensuring that features of the same age are closely clustered, while features of different ages are contrasted in alignment with age progression. To further improve this process, we introduce a soft proxy matching loss, which assigns weights to negative pairs based on their absolute age differences, thereby modeling the relationships between features and proxies proportionally. This loss function pushes negative pairs with greater age differences farther apart than those with smaller differences. As illustrated in Fig. 1, the top plot shows learned age features mapped in alignment with age progression, while the bottom plot demonstrates how features from the same individual at different ages are positioned in the direction of age progression.
@@
The proposed method is versatile. It can be directly applied to age estimation by learning an age feature space that aligns facial features along the direction of aging. For AIFR, the method builds upon our preliminary work [25] by simultaneously learning both age and identity features. To ensure that identity features are robust to age variations, the gradient for age information is gradually reversed using a gradient reversal layer (GRL) [26]. The key distinction between our previous work [25] and the proposed method lies in the explicit modeling of age progression in OrdCon, which is critical for both age estimation and AIFR tasks.
@@
  • We introduce Order-Enhanced Contrastive Learning (OrdCon) for robust and generalized age feature extraction. Unlike conventional contrastive learning methods that contrast image pairs, OrdCon contrasts direction vectors between feature pairs along the aging direction, as well as age features with age-specific proxies.
  • We propose a soft proxy matching loss that assigns weights to negative pairs in the metric learning objective based on absolute age differences, allowing for more nuanced modeling of age progression while reducing intra-class variance.
  • The proposed method is versatile, as it can be applied to both age estimation and AIFR tasks.
  • We provide empirical evidence that OrdCon achieves state-of-the-art or comparable performance in homogeneous-dataset experiments, and superior performance in cross-dataset settings, demonstrating its robustness and generalization capabilities.
The rest of this paper is organized as follows: Section II reviews related work on age estimation and AIFR, focusing on feature extraction methods. In Section III, we provide a detailed description of OrdCon, including problem formulation, the definition of contrastive objectives, and its adaptation for age estimation and AIFR. Section IV covers the experimental settings and presents the performance of OrdCon compared to state-of-the-art methods in both homogeneous and cross-dataset experiments. Finally, Section V concludes our work.
@@ II Related Work
In this section, we review existing works on age estimation and AIFR, focusing on age feature extraction techniques used in both tasks. We also discuss the learning techniques commonly employed for these tasks.
II-A Age Estimation
@@
Over the past few decades, various methods for face-based age estimation have been developed. Early approaches, like Kwon and Lobo’s work [27], classified faces into three age groups using craniofacial development theory and wrinkle analysis. Similarly, Wu et al. [28] highlighted the importance of wrinkles in modeling facial aging and estimating age.
Before the rise of deep learning [29], handcrafted feature extraction techniques were the primary methods used for age estimation. Edwards et al. [30] developed the Statistical Face Model, which Lanitis et al. [31] used to extract age-related facial features and correlate them with age labels. Geng et al. [32] introduced the AGing pattErn Subspace (AGES), which projected unseen faces into a subspace based on learned aging patterns. Gao et al. [33] proposed Fuzzy Linear Discriminant Analysis (LDA) to address the ambiguity of mapping ages to age groups, using an Age Membership Function that incorporated age relevance into LDA classification. Similarly, Guo et al. [34] introduced a kernel-based regression method for age estimation. Another prominent feature extraction approach from this period is the Biologically Inspired Features (BIF) [1], which uses Gabor filters followed by max-pooling, proving effective for age estimation.
@@
With the advent of larger age-labeled datasets [35, 36], convolutional neural networks (CNNs) became the dominant feature extraction method. One of the early CNN-based approaches was [37] in which a two-layer CNN was used for age prediction. Han et al. [38] later proposed a modified version of AlexNet [39] for multi-task learning, predicting age along with other facial attributes simultaneously.
CNN-based age estimation methods can be broadly categorized into two groups, both focusing on enhancing feature extraction. The first group achieves this by designing customized network architectures. For example, Yi et al. [40] proposed a multi-stream CNN, where multiple facial patches containing high-dimensional information are processed independently before being fused at the output layer. Angeloni et al. [41] and Chen et al. [42] also used multi-stream strategies, focusing on key facial attributes such as the eyes, nose, and mouth. Wang et al. [43] leveraged multiple facial patches learned by an Adaboost algorithm, which were sequentially fed into a customized CNN based on how much age information they carry. More recently, Xia et al. [13] enhanced feature extraction by introducing a multi-stage feature constraints learning framework, progressively refining features through three stages to align them with the corresponding age range.
@@
The second group of methods aims to improve age feature extraction by effectively modeling the relationship between data and labels through learning objectives. For example, Niu et al. [44] framed age estimation as an ordinal regression problem, using a parallel set of fully connected layers to handle binary classification sub-problems for each age. Chen et al. [45] introduced Ranking-CNN to consider ordinal relationships between ages, while Pan et al. [46] proposed a mean-variance loss to concentrate age predictions around the ground truth. Shen et al. [47] tackled inhomogeneity in the mapping between facial features and age by attaching deep forests to CNNs, allowing random forests to perform soft data partitioning and learn age distributions.
At the same time, Lim et al. [21] introduced an order learning framework that compares instances in pairs to establish their relative order rather than predicting an absolute age. Later, Lee et al. [48] integrated order learning with metric learning to create a feature space in which distances between instances reflect both the order and magnitude of age differences.
Recently, there has been growing interest in using contrastive learning for age estimation. Pan et al. [49] revised the contrastive loss to leverage pairwise ordinal relationships between images, ensuring that the distance between features is proportional to the age difference. Li et al. [50] introduced a group-aware contrastive network that minimizes intra-group variance and maximizes inter-group distances by selecting positive and negative sample pairs based on age groups. Most recently, Zha et al. [51] proposed a contrastive learning framework that simultaneously ranks and contrasts features to maintain the continuity and smooth progression of aging.
Unlike the previous order learning and contrastive learning methods, the proposed approach combines the strengths of both while introducing a soft proxy matching loss for metric learning to provide a more structured representation of aging. Compared to [21] and [48], OrdCon learns a generalized age feature space by adopting contrastive learning concepts to contrast direction vectors with age progression and features with age-specific proxies. Compared to existing contrastive learning-based methods, OrdCon not only considers age differences between samples but also maps the age feature in the direction that aligns with the age progression to effectively organize features in the feature space. Additionally, OrdCon demonstrates robustness and generalization in both age estimation and AIFR tasks, whereas prior methods were designed for only one.
II-B Age-Invariant Face Recognition
An early approach to AIFR was introduced by Park et al. [11], who used 3D modeling to simulate facial aging and mitigate age-related variations in face recognition. This method generated aged faces from 2D images by constructing 3D models incorporating both shape and texture, though the results were limited by the accuracy of age estimation algorithms. Later, Li et al. [7] categorized AIFR methods into two types: generative and discriminative. Generative models, like Park et al.’s, synthesize faces matching the target age before recognition, whereas discriminative approaches directly extract age-invariant features for recognition.
Due to the low quality of samples synthesized by early generative models, most existing AIFR methods are discriminative [52, 53, 54, 55, 9]. Studies on human aging [56, 57] indicate that facial aging is associated with skin texture, which becomes rougher over time. Consequently, early discriminative approaches [7, 58, 59, 54] employed local binary pattern (LBP) techniques to extract facial features, followed by dimensionality reduction methods like principal component analysis (PCA) or LDA. Gong et al. [53] modeled the extracted low-dimensional features as a combination of multiple components, with one component representing age information, which could be separated from the global features before performing recognition.
@@ II-B Age-Invariant Face Recognition
With the increasing popularity of CNNs, researchers have begun using them as feature extractors in discriminative AIFR methods. Wen et al. [55] were the first to use a CNN for AIFR, designing a customized network with latent identity analysis to learn disentangled features. Zheng et al. [60] proposed a multi-task framework with one path for face recognition and another for age estimation, directly subtracting age features from global features to obtain age-invariant identity features. Later, Wang et al. [61] adopted a similar multi-task strategy, introducing a novel decomposition method to separate age features from identity features using a spherical coordinate system. They also applied a regression loss to learn finer age features, enhancing the decomposition process. Subsequently, Wang et al. [62] proposed a discriminative method based on adversarial learning and a canonical mapping module to reduce the correlation between age and identity features. This adversarial learning approach demonstrated superior performance on several benchmark datasets compared to previous methods.
With the significant improvement in synthesized image quality, researchers have shifted their focus back to generative approaches. Zhao et al. [63] proposed an end-to-end method that simultaneously synthesizes faces across different age groups and performs feature disentanglement using a GRL to reverse the gradient of age information during back-propagation. Huang et al. [64] presented a similar framework, benefiting from more realistic faces generated by their synthesis model. Zhao et al. [65] further proposed a Generative Adversarial Network (GAN) [66] for age-invariant face verification (AIFV), synthesizing realistic facial images for different age groups by manipulating latent features between the encoder and decoder. Recently, the Multi-Feature Fusion and Decomposition framework [67] was introduced to enhance feature extraction by combining facial features from images using multi-head self-attention [68], along with a decomposition approach to separate age and identity features.
@@
The aforementioned methods rely on conventional supervised training, directly mapping features to labels without explicit representation learning. However, contrastive learning has been largely unexplored for AIFR. The notable exception is Wang et al. [25], which maximized similarity between features from different age groups of the same identity using additional samples from a face synthesis model. However, this approach did not explicitly model the continuous progression of aging, which may lead to inaccurate feature representations due to entanglement between age and identity features.
@@
In this work, we focus on generalized age feature extraction by initially learning age and identity features simultaneously. Once natural age progression is learned through various contrastive objectives, the age-related variation is gradually minimized to obtain age-invariant identity features.
@@
Figure 2: Workflow of the proposed method for age estimation and AIFR, compared to our preliminary work [25]. For age estimation, OrdCon processes an image pair with age differences to learn the ordinal relationship between images. For AIFR, CACon employs a GAN model to synthesize a third sample in a different age group and minimizes the distances among samples from various age groups. Extending this approach, OrdCon synthesizes samples for each predefined age group to learn the individual’s aging process in a multitask framework. To achieve age-invariant identity features, the gradient of age information is gradually reversed using a GRL.

### v2

@@ Abstract
Generalized age feature extraction is crucial for age-related facial analysis tasks, such as age estimation and age-invariant face recognition (AIFR). Despite the recent successes of models in homogeneous-dataset experiments, their performance drops significantly in cross-dataset evaluations. Most of these models fail to extract generalized age features as they only attempt to map extracted features with training age labels directly without explicitly modeling the natural ordinal progression of aging. In this paper, we propose Order-Enhanced Contrastive Learning (OrdCon), a novel contrastive learning framework designed explicitly for ordinal attributes like age. Specifically, to extract generalized features, OrdCon aligns the direction vector of two features with either the natural aging direction or its reverse to model the ordinal process of aging. To further enhance generalizability, OrdCon leverages a novel soft proxy matching loss as a second contrastive objective, ensuring that features are positioned around the center of each age cluster with minimal intra-class variance and proportionally away from other clusters. By modeling the ageing process, the framework can enhance generalizability by improving the alignment of samples from the same class and reducing the divergence of direction vectors. We demonstrate that our proposed method achieves comparable results to state-of-the-art methods on various benchmark datasets in homogeneous-dataset evaluations for both age estimation and AIFR. In cross-dataset experiments, OrdCon outperforms other methods by reducing the mean absolute error by approximately 1.38 on average for the age estimation task and boosts the average accuracy for AIFR by 1.87%.
@@ I Introduction
Despite the importance of generalized age features, current age feature extraction methods mainly rely on conventional supervised learning, where models are designed to capture facial features associated with aging by directly mapping them to training labels [13], learning particular features tied to the age distributions and demographic characteristics in the training data [14]. A common approach involves using deep neural networks to automatically learn age features, such as skin texture, wrinkles, and changes in facial structure [15, 16]. While these methods have demonstrated success in controlled datasets and environments, they fall short of generalizing to diverse scenarios with unfamiliar variability in aging patterns across individuals [17, 18], yielding poor alignment of samples from the same class on unseen data [19].
@@
Additionally, models trained with conventional supervised learning techniques often struggle to capture the inherent continuity and progression of age-related changes [20]. Most supervised approaches treat each sample independently, disregarding the ordinal nature of the aging process [14, 15]. This inability to capture smooth transitions in facial features over time can result in inconsistent or unrealistic predictions that do not reflect the continuous nature of aging [21].
@@
Figure 1: Visual representation of the age features learned using the proposed method. Top: Feature space aligned with the direction of age progression. Bottom: Features from two individuals at different ages, following the learned age progression. Best viewed in color.
@@
To address the aforementioned limitations, we propose a new contrastive learning framework called Order-Enhanced Contrastive Learning (OrdCon). OrdCon enhances the generalizability of age feature extraction in two ways. First, it leverages order learning [22] to model how features evolve with age progression explicitly. Unlike other contrastive learning frameworks that contrast either image pairs [23, 24] or absolute semantic differences between samples [25], OrdCon contrasts the direction vector between features in the feature space by aligning the direction vectors from younger to older faces with the natural direction of age progression. This approach enables the model to capture continuous and gradual transitions in facial features, effectively addressing the limitation of conventional methods in modeling progression.
@@
Secondly, to complement the directional alignment of features along the aging progression and further boost the generalizability, OrdCon leverages a novel soft proxy matching loss as the second contractive objective to reduce intra-class variance and create a structured feature space. In this approach, proxies are learnable entities that represent the centers of specific age clusters. This loss contrasts extracted features relative to these proxies, effectively minimizing intra-class variance by ensuring that features of the same age are closely clustered. Moreover, it assigns weights to negative pairs based on their absolute age differences, thereby pushing negative pairs with greater age differences farther apart than those with smaller differences to model the relationships between features and proxies proportionally. As illustrated in Fig. 1, the top plot shows learned age features mapped in alignment with age progression, while the bottom plot demonstrates how features from the same individual at different ages are positioned in the direction of age progression.
@@
The resulting framework for extracting generalized age features is inherently versatile. It can be directly applied to age estimation by learning an age feature space that aligns facial features along the direction of aging. For AIFR, the method builds upon our preliminary work [26], Cross-Age Contrastive Learning (CACon), by simultaneously learning both age and identity features. To ensure that identity features are robust to age variations, the gradient for age information is gradually reversed using a gradient reversal layer (GRL) [27]. The key distinction between OrdCon and CACon lies in the explicit modeling of age progression, which is critical for improving the generalizability of both age estimation and AIFR tasks.
Since this paper focuses on applications, the theoretical analysis of the generalizability of contrastive learning for regression problems is out of scope and could be written as a separate paper. To demonstrate the generalizability of our framework, in addition to its performance in cross-dataset experiments, we adapt the matrix alignment of positive samples proposed in [19] to measure the similarity of augmented samples of the same class after contrastive learning. In addition, inspired by another matrix, the divergence of class centers proposed in [19], we propose a new matrix called the divergence of direction vectors to better align with the objective of regression problems.
@@
  • We introduce a new contrastive learning framework called Order-Enhanced Contrastive Learning (OrdCon) for robust and generalized age feature extraction. Unlike conventional contrastive learning methods that contrast image pairs, OrdCon contrasts direction vectors between feature pairs along the aging direction, as well as age features with age-specific proxies.
  • We propose a soft proxy matching loss that assigns weights to negative pairs based on absolute age differences, thereby allowing for reduced intra-class variance and proportionally separated clusters for improved generalizability.
  • To the best of our knowledge, this paper is the first to study and quantify the generalizability for models dealing with ordinal data.
  • We demonstrate that the proposed method is versatile, as it can be applied to both age estimation and AIFR tasks.
  • We provide empirical evidence that OrdCon achieves state-of-the-art or comparable performance in homogeneous-dataset experiments, and superior performance in cross-dataset settings.
The rest of this paper is organized as follows: Section II reviews related work on contrastive learning and order learning techniques, followed by a discussion of related works on age estimation and AIFR. In Section III, we provide a detailed description of OrdCon, including problem formulation, the definition of contrastive objectives, and its adaptation for age estimation and AIFR. Section IV presents the experimental settings and compares the performance of OrdCon to state-of-the-art methods in both homogeneous and cross-dataset experiments. Section V discusses the applicability and limitations of OrdCon, and Section VI concludes our work.
@@ II Related Work
Given that the primary focus of OrdCon is feature extraction, this section reviews key works in two core feature extraction techniques, contrastive learning and order learning, upon which our work is built. Then, we review domain-specific methods on age estimation and AIFR.
II-A Contrastive Learning frameworks
Contrastive learning has emerged as a dominant and highly effective paradigm for representation learning. Among early contributions, SimCLR [23] demonstrated its effectiveness by learning a feature space where representations of positive pairs (i.e., semantically similar samples) are pulled closer together. In contrast, representations of negative pairs (i.e., dissimilar samples) are pushed far apart. SimCLR requires a large batch size for training. Its samples are generated using a composite data augmentation workflow where positive pairs originate from the same input image, while negative pairs are derived from different samples, even if they share the same class label.
Other seminal unsupervised contrastive learning frameworks include the Momentum Contrast (MoCo) family [28, 29, 30], Bootstrap Your Own Latent (BYOL) [31], and SimSiam [32]. The MoCo family aims to address the practical challenge of requiring huge batch sizes for training contrastive learning frameworks. It achieves this by maintaining a dynamic dictionary of encoded representations from preceding mini-batches. This dictionary is updated with features from a slowly progressing, momentum-updated encoder, thereby providing a significant and consistent set of negative examples without the computational overhead of a massive batch. Conversely, BYOL and SimSiam attempt to alleviate the representation collapse issue by leveraging encoders with asymmetric architectures to prevent direct parameter copying.
While the aforementioned methods are designed for unlabeled data, Khosla et al. [24] proposed Supervised Contrastive Learning (SupCon) to leverage label information within contrastive learning. The objective then becomes to pull together the representations of all samples within a class cluster, while simultaneously pushing this cluster away from all other class clusters. SupCon was shown to consistently outperform standard cross-entropy loss, particularly in terms of representation quality.
The superior generalizability of contrastive learning frameworks has garnered increasing attention from researchers. Huang et al. [19] reveal three factors that contribute to that. They are the alignment of positive samples, divergence of class centers, and concentration of augmented data. The first two are related to learning objectives, and the data augmentation process determines the last one. Recently, Hieu and Ledent [33] proposed a generalization analysis for supervised contrastive representation learning in the practical, non-i.i.d. setting, where a fixed pool of labeled data is recycled to form training tuples. They derive excess risk bounds using a novel theoretical framework inspired by U-statistics [34].
Although contrastive learning has proven transformative for classification tasks, it is not directly applicable to tasks with continuous target variables. The main challenge lies in defining positive and negative pairs for regression tasks.
To this end, Li et al. [35] introduced a group-aware contrastive network that minimizes intra-group variance and maximizes inter-group distances by selecting positive and negative sample pairs based on age groups. Similarly, Pitawela et al. [36] use samples with the same label to construct positive samples and samples with different labels to construct negative samples. They also introduced a configurable hyperparameter to determine the distance among clusters in the feature space. Conversely, Zha et al. [25] define positive and negative pairs based on absolute age differences among samples. Specifically, for a given anchor sample, any other sample in the batch can form a positive pair, while the corresponding negative pairs are all samples whose label distance to the anchor is greater than that of the positive pair.
Compared to the aforementioned contrastive learning frameworks for regression tasks, OrdCon not only considers absolute age differences between samples but also leverages ordinal information between samples by contrasting direction vectors. This ensures that learned features are mapped in a direction that aligns with age progression, resulting in better generalizability.
II-B Order learning
The idea of order learning was initially proposed by Lim et al [22]. The authors introduced a new approach to problems with inherently ordered labels by focusing on relative comparisons rather than absolute value estimation. Instead of directly predicting a continuous value or performing multiple binary classifications, their method reframed the problem as determining the relative order between two instances. Later, Lee et al. [37] integrated order learning with metric learning to create a feature space in which distances between instances reflect both the order and magnitude of age differences. Recently, Unsupervised Order Learning was proposed, which integrates the idea of order learning into clustering for ordinal data [38].
Different from the above order learning methods, OrdCon focuses on learning a generalized feature space by adopting contrastive learning to contrast direction vectors with age progression and features with age-specific proxies. Moreover, we tested its versatility on two age-related tasks rather than only one, like previous methods.
II-C Age Estimation
@@
Over the past few decades, various methods for face-based age estimation have been developed. Early approaches, like Kwon and Lobo’s work [39], classified faces into three age groups using craniofacial development theory and wrinkle analysis.
@@
With the advent of larger age-labeled datasets [40, 41], convolutional neural networks (CNNs) became the dominant feature extraction method. One of the early CNN-based approaches was [42] in which a two-layer CNN was used for age prediction.
@@
A primary challenge in age estimation is that age is not a discrete, independent class but an ordinal, continuous variable. To this end, Niu et al. [43] and Chen et al. [44] framed age estimation as an ordinal regression problem, using a parallel set of fully connected layers to handle binary classification sub-problems for each age. Pan et al. [45] proposed a mean-variance loss to concentrate age predictions around the ground truth. Shen et al. [46] tackled inhomogeneity in the mapping between facial features and age by attaching deep forests to CNNs, allowing random forests to perform soft data partitioning and learn age distributions.
To mine the ordinal relationships among ages, label distribution learning (LDL) [47] as been widely adapted by the research community. In LDL, instead of using a single age label, it represents the ground truth as a soft probability distribution over a range of ages. This allows a single training sample to contribute to the learning of its chronological age, as well as the ages of its neighbors. Recent advances in LDL have focused on designing sophisticated loss functions. Specifically, Adaptive LDL [48, 49] has been proposed to generate instance-aware, unimodal distributions reflecting sample-specific ambiguity. On the other hand, ordinal LDL [50] aims to explicitly models the sequential and semantic ordinality of age by incorporating spatial, semantic, and temporal order relationships into its loss function. Although LDL-based methods works well on standard benchmarks, they perform poorly on cross-dataset evaluations [51].
Unlike the aforementioned age estimation methods, our approach primarily focuses on improving the generalizability of age estimation models by combining contrastive learning and order learning techniques. In addition, the proposed framework has been demonstrated to be able to tackle both age estimation and AIFR problems rather than just a single usage.
II-D Age-Invariant Face Recognition
@@
The mainstream strategy in modern AIFR is feature disentanglement, which aims to separate identity-specific features from age-related characteristics. Zheng et al. [52] proposed a multi-task framework with one path for face recognition and another for age estimation, directly subtracting age features from global features to obtain age-invariant identity features. Wang et al. [53] adopted a similar multi-task strategy, introducing a novel decomposition method to separate age features from identity features using a spherical coordinate system. They also applied a regression loss to learn finer age features, enhancing the decomposition process. Recently, Xie et al. [54] proposed an Implicit and Explicit Feature Purification (IEFP) framework that uses a regularizer to minimize the mutual information between the identity-focused features and features from a pre-trained age-estimation network.
Another strategy gaining popularity is combining generative and discriminative models to form unified frameworks. The Age-Invariant Model (AIM) [10] was one of the first to perform cross-age face synthesis and recognition simultaneously. Recently, Yao et al. [55] evaluated multiple age synthesis models and demonstrated that using synthesized faces at different ages can enhance the AIFR model’s performance.
@@
Like existing age estimation models, the above AIFR models are all designed for homogeneous-dataset experiments. Techniques to improve generalizability, such as contrastive learning frameworks, have been largely unexplored for AIFR. The only notable work is our preliminary work, CACon [26], which maximized similarity between features from different age groups of the same identity using additional samples from a face synthesis model. However, this approach did not explicitly model the continuous progression of aging, which may lead to inaccurate feature representations due to entanglement between age and identity features.
@@
In this work, we focus on generalized age feature extraction by initially learning age and identity features simultaneously. Once natural age progression is learned through the two contrastive objectives, the age-related variation is gradually minimized to obtain age-invariant identity features.
@@
Figure 2: Proposed framework for age estimation and AIFR, compared to our preliminary work [26]. For age estimation, OrdCon processes an image pair with age differences to learn the ordinal relationship between images. For AIFR, CACon uses a third sample from a different age group to minimize the distances among samples from various age groups. Extending this approach, OrdCon leverages samples for each predefined age group to learn the individual’s aging process in a multitask manner. To achieve age-invariant identity features, the gradient of age information is gradually reversed using a GRL.